According to the following reaction, how many grams of ammonium chloride are needed to form 30.8 grams of hydrochloric acid?

ammonium chloride (aq) ammonia (g) + hydrochloric acid (aq)
1
Expert's answer
2018-02-12T07:32:39-0500
. m (HCl) =30.8 g
.m (NH4Cl)=?
The reaction is
NH4Cl= NH3 + HCl
Let's find the amount of substance (HCl)= m(HCl)/ Mr(HCl)
Mr (HCl)= Ar(H)+ Ar(Cl)=1+35.5= 36.5
. n=30.8/36.5=0.84 mol
. n (NH4Cl) = n(HCl)=0.84 mol
.m (NH4Cl)= Mr(NH4Cl) *n (NH4Cl)
Mr(NH4Cl)= Ar(N)+ 4 Ar(H)+ Ar (Cl)= 14+4+35.5=53.5
m (NH4Cl)=53.5 *0.84=44.94 g